OpenStudy (anonymous):

Solve: (x - 8)(x + 11) = 0. A: {8, -11} B: {-8, 11} C: {0, 11/8} D: no solution

OpenStudy (anonymous):

This should have two solutions....right?

OpenStudy (anonymous):

the question is just A B C or D it doesnt have an option to pick more than one

OpenStudy (anonymous):

But one of the answers given is one of the solutions. Anyways...what you want to do is pick a number for x in the first parentheses to make the inside of the parentheses zero.

OpenStudy (anonymous):

(x-8) what number can we pick for x to make the inside zero?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Great that's pretty much all you have to do. Find the answer where x=8 or in other words (8,?)

OpenStudy (anonymous):

wow that wasnt that hard

OpenStudy (anonymous):

I put a ? because we don't know what y is and we don't have to exactly because we know that x= 8

OpenStudy (anonymous):

Well this is a special case. So we can't always use this method

OpenStudy (anonymous):

Solve: (x - 8)(x + 11) = 0. (Points : 1) {8, -11} {-8, 11} {0, 11/8} no solution

OpenStudy (anonymous):

Because the left side equals zero. And we have two parentheses multiplying we can easily find the answer by making the insides of the paraentheses zero

OpenStudy (anonymous):

Solve the equation (4x + 1)(3x - 2) = 0 (Points : 1) {-4, 1.5} {-1/4, 2/3} {1/4, 2/3} {-4, 3}

OpenStudy (anonymous):

what about this

OpenStudy (anonymous):

same method?

OpenStudy (anonymous):

well look at it this way we want the inside of a parentheses to be zero (4x + 1)(3x - 2) Because if we have (0)*(3x - 2) =0 since anything we multiply with zero is zero. And we want the left side to be equal to zero.

OpenStudy (anonymous):

So yeah use the same method.

OpenStudy (anonymous):

We want 4x + 1=0

OpenStudy (anonymous):

Now solve for x

OpenStudy (anonymous):

ok thanks